# On semipositone problems over ${\mathbb{R}}^{N}$ for the fractional $p$-Laplace operator

### v3

@@ 3. Existence and qualitative properties of the mountain pass solutions
Uniform bound of the positive part: For $M>0$, define $(u_{a}^{+})_{M}=\min\{(u_{a})^{+},M\}$ for every $a\in(0,a_{1})$. Clearly $(u_{a}^{+})_{M}\geq 0$ and $(u_{a}^{+})_{M}\in L^{\infty}({\mathbb{R}}^{N})\cap{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Fixed $\sigma\geq 1$, define $\phi=((u_{a}^{+})_{M})^{\sigma}$. Then, $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}.$ By Theorem 3.1, $u_{a}$ is a weak solution of (SP). Taking $\phi$ as a test function, we write
@@
Using ($\bf i{2}$), Remark 3.2, and ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, we have the following lower bound of the L.H.S:
@@
Step 1: In this step, for $\sigma=p^{*}_{s}$, we show that the set $\{(u_{a}^{+})^{\sigma+p-1}:a\in(0,a_{1})\}$ is uniformly bounded on $L^{\frac{\sigma}{p}}({\mathbb{R}}^{N})$. From ($\bf\tilde{f{1}}$), for any $\epsilon>0$, there exists $C=C({\epsilon},a_{1})$ such that $\lvert f_{a}(u_{a})\rvert\leq C+\epsilon\lvert u_{a}\rvert^{p^{*}_{s}-1}$ for all $a\in(0,a_{1})$. Hence for all $a\in(0,a_{1})$, we also have
$\lvert f_{a}(u_{a})\rvert\leq C+\epsilon(u_{a}^{+}+u_{a}^{-})^{p^{*}_{s}-1}=C+\epsilon\left((u_{a}^{+})^{p^{*}_{s}-1}+(u_{a}^{-})^{p^{*}_{s}-1}\right).$
Using the above inequality and noting that $u_{a}^{+}$ and $u_{a}^{-}$ have disjoint supports, we estimate the R.H.S of (3.8) as follows:
$\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\leq C\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x+{\epsilon}\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x,$
where using ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, the uniform boundedness of $\{u_{a}:a\in(0,a_{1})\}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ (Theorem 3.1), and Remark 3.2,
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\leq C(N,s,p)\left\lVert g\right\rVert_{\infty}\left\lVert u_{a}^{+}\right\rVert_{s,p}^{p^{*}_{s}}\leq C(N,s,p,f,g,a_{1}),\quad\forall\,a\in(0,a_{1}),$
and by Hölder’s inequality with the conjugate pair $(\frac{p^{*}_{s}}{p^{*}_{s}-p},\frac{p^{*}_{s}}{p})$, we get
@@
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x\leq C(N,s,p,f,g,a_{1})\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}.$
Thus for any ${\epsilon}>0$, from (3.8) we have
$\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}$ $\leq\left(\frac{p^{*}_{s}+p-1}{p}\right)^{p}$
$\left(C_{1}+C_{2}{\epsilon}\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\right),$ (3.9)
where $C_{1}=C_{1}(N,s,p,{\epsilon},f,g,a_{1})$ and $C_{2}=C_{2}(N,s,p,f,g,a_{1})$. Choose ${\epsilon}>0$ so that
@@
Then (3) implies
@@
$\delta_{0}\leq\frac{\|u_{a}\|_{s,p}^{p}}{p}\leq\frac{1}{p}\int_{{\mathbb{R}}^{N}}g(x)|f_{a}(u_{a})u_{a}|\,{\rm d}x\leq C(f,a_{2})\|g\|_{1}(\|u_{a}\|_{\infty}+\|u_{a}\|_{\infty}^{p_{s}^{*}}).$ (3.16)
@@
Now we define $w_{n}=u_{n}-\tilde{u}$. Since $u_{n},\tilde{u}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{r}({\mathbb{R}}^{N})$ (Proposition 3.3 and Proposition 3.5), we have $w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty]$. Hence $\lvert w_{n}\rvert^{q-1}w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ for every $q\geq 1$.
For $p\geq 2$: We use [30, Lemma 2.3] to obtain the following lower bound of the R.H.S of (3):
@@
where $C=C(N,s,p,f,a_{2})$ and $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Define $\gamma:=\frac{p^{*}_{s}}{p\bar{r}^{\prime}}.$ Using the definition of $\bar{r}$ in (3.23) it follows that $\gamma>1$. We consider two sequences $\{l_{j}\}$ and $\{m_{j}\}$ by the following recursive process:
$l_{0}=p_{s}^{*},\quad l_{j+1}=\gamma l_{j}+\frac{p^{*}_{s}(p-1)}{p},\quad m_{j}=\frac{l_{j}}{\bar{r}^{\prime}}.$
@@
$\leq\left(\widetilde{C}m_{j}^{p}\right)^{\gamma\bar{r}^{\prime}}\left\lVert g\right\rVert_{\infty}^{\gamma}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\gamma}$
$\leq\left(\widetilde{C}m_{j}^{p}\right)^{\gamma\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\gamma},$ (3.26)
for some $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. By Lemma A.2 we observe that $m_{j}\sim\frac{\gamma^{j}}{\bar{r}^{\prime}}$ as $j\rightarrow\infty$. Thus iterating on (3) we obtain
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x$ $\leq\prod_{i=0}^{j-1}\left(\widetilde{C}m_{i}^{p}\right)^{\gamma^{j-i}\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{0}}\,{\rm d}x\right)^{\gamma^{j}}$
$\leq\left(\widetilde{C}\right)^{\sum\limits_{i=0}^{j-1}\gamma^{j-i}\bar{r}^{\prime}}\gamma^{p\bar{r}^{\prime}\sum\limits_{i=0}^{j-1}i\gamma^{j-i}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\gamma^{j}}.$ (3.27)
Notice that $S_{1}:=\sum\limits_{i=0}^{\infty}\gamma^{-i}<\infty$ and $S_{2}:=\sum\limits_{i=0}^{\infty}i\gamma^{-i}<\infty$. Take $C>\max\{1,\widetilde{C}\}$. Then for all $n\geq n_{1},j\in{\mathbb{N}}$ using (3) we deduce
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\leq C^{\gamma^{j}S_{1}}\gamma^{p\bar{r}^{\prime}\gamma^{j}S_{2}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\gamma^{j}}.$
@@
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C^{\frac{\gamma^{j}}{l_{j}}S_{1}}\gamma^{p\bar{r}^{\prime}\frac{\gamma^{j}}{l_{j}}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\frac{\gamma^{j}}{l_{j}}}\leq C^{\beta_{2}S_{1}}\gamma^{p\bar{r}^{\prime}\beta_{2}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\beta_{1}},$
where $\beta_{1},\beta_{2}>0$ have been chosen such that for all $j\in{\mathbb{N}}$, $\beta_{1}<\frac{\gamma^{j}}{l_{j}}<\beta_{2}$ (see Lemma A.2). Therefore, there exists $C>1$ such that for all $n\geq\max\{n_{1},n_{2}\}$ and $j\in{\mathbb{N}}$ large enough
@@
For $\frac{2N}{N+2s}<p<2$: For any $q\geq 1$, we take $\lvert w_{n}\rvert^{q-1}w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ as a test function in (3) and use [30, Lemma 2.4] to obtain
@@
where $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Define $\gamma:=\frac{p^{*}_{s}}{2\bar{r}^{\prime}}.$ Using the definition of $\bar{r}$ in (3.30) we can verify that $\gamma>1$. We consider two sequences $\{l_{j}\}$ and $\{m_{j}\}$ by the following recursive process:
$l_{0}=p_{s}^{*},\quad l_{j+1}=\gamma l_{j}+\frac{p^{*}_{s}}{2},\quad m_{j}=\frac{l_{j}}{\bar{r}^{\prime}}.$
@@
$\leq\left(\widetilde{C}m_{j}^{2}\right)^{\gamma\bar{r}^{\prime}}\left\lVert g\right\rVert_{\infty}^{\gamma}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\gamma}$
$\leq\left(\widetilde{C}m_{j}^{2}\right)^{\gamma\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\gamma},$ (3.32)
@@
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\leq C^{\gamma^{j}S_{1}}\gamma^{2\bar{r}^{\prime}\gamma^{j}S_{2}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\gamma^{j}},$
@@
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C^{\frac{\gamma^{j}}{l_{j}}S_{1}}\gamma^{2\bar{r}^{\prime}\frac{\gamma^{j}}{l_{j}}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\frac{\gamma^{j}}{l_{j}}}\leq C^{\alpha_{2}S_{1}}\gamma^{2\bar{r}^{\prime}\alpha_{2}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\alpha_{1}},$
where $\alpha_{1},\alpha_{2}$ have been chosen such that for all $j\in{\mathbb{N}}$, $\alpha_{1}<\frac{\gamma^{j}}{l_{j}}<\alpha_{2}$ (see Lemma A.2). Finally, we find a $C>1$ such that for all $n\geq\max\{n_{1},n_{3}\}$ and $j\in{\mathbb{N}}$ large enough
@@
$\Psi(\varrho)=\mathcal{H}^{N-2}(\mathbb{S}^{N-1})\int_{-1}^{1}\frac{(1-t^{2})^{\frac{N-3}{2}}}{(1-2t\varrho+\varrho^{2})^{\frac{N+sp}{2}}}\,\mathrm{d}t.$ (3.37)
@@
$|x-y|~{}\geq~{}|x|-|y|~{}\geq~{}|x|-\varepsilon~{}\geq~{}|x|-\frac{|x|}{2}=\frac{|x|}{2}.$
@@
where $C_{1}(\beta,\varepsilon)<0$. Since $a_{n}\rightarrow 0$, there exists $n_{1}\in{\mathbb{N}}$ such that $a_{n}B\leq-C_{1}(\beta,\varepsilon){\varepsilon}^{(\beta+1)(p-1)}$ and $u_{a_{n}}\geq 0$ a.e. in ${\mathbb{R}}^{N}$ (Proposition 3.8) for all $n\geq n_{1}$. Thus, for all $n\geq n_{1}$ using the assumption (3.38), following holds weakly in $\overline{B_{R_{1}}}^{c}$:

### v4

@@ 3. Existence and qualitative properties of the mountain pass solutions
Uniform bound of the positive part: For $M>0$, define $(u_{a}^{+})_{M}=\min\{(u_{a})^{+},M\}$ for every $a\in(0,a_{1})$. Clearly $(u_{a}^{+})_{M}\geq 0$ and $(u_{a}^{+})_{M}\in L^{\infty}({\mathbb{R}}^{N})\cap{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Fixed $\sigma\geq 1$, define $\phi=((u_{a}^{+})_{M})^{\sigma}$. Now we claim that $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. First, we recall the following inequality from [30, (2.4), Page 1359]: for any $\alpha,\beta\in\mathbb{R}$ and $\sigma\geq 1$, we have
$\left||\alpha|^{\sigma-1}\alpha-|\beta|^{\sigma-1}\beta\right|\leq\sigma\left(|\alpha|^{\sigma-1}+|\beta|^{\sigma-1}\right)|\alpha-\beta|.$ (A)
Taking $\alpha=(u_{a}^{+})_{M}(x)$ and $\beta=(u_{a}^{+})_{M}(y)$ into (A) and using the fact that $0\leq(u_{a}^{+})_{M}\in L^{\infty}({\mathbb{R}}^{N})$, we get
$\left|\phi(x)-\phi(y)\right|$ $\leq\sigma\left(|(u_{a}^{+})_{M}(x)|^{\sigma-1}+|(u_{a}^{+})_{M}(y)|^{\sigma-1}\right)\left|(u_{a}^{+})_{M}(x)-(u_{a}^{+})_{M}(y)\right|$
$\leq 2\sigma\|(u_{a}^{+})_{M}\|_{\infty}^{\sigma-1}\left|(u_{a}^{+})_{M}(x)-(u_{a}^{+})_{M}(y)\right|.$
Since $(u_{a}^{+})_{M}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we have $\left\lVert(u_{a}^{+})_{M}\right\rVert_{s,p}<\infty$ and which further implies $\left\lVert\phi\right\rVert_{s,p}<\infty$ i.e.,
$\left\lVert\phi\right\rVert_{s,p}^{p}=$ $\iint_{{\mathbb{R}}^{2N}}\frac{\lvert\phi(x)-\phi(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}\leq\left(2\sigma\|(u_{a}^{+})_{M}\|_{\infty}^{\sigma-1}\right)^{p}\iint_{{\mathbb{R}}^{2N}}\frac{\lvert(u_{a}^{+})_{M}(x)-(u_{a}^{+})_{M}(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}<\infty.$
Also, $(u_{a}^{+})_{M}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{\infty}({\mathbb{R}}^{N})$ implies $(u_{a}^{+})_{M}\in L^{p_{s}^{*}}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, which further implies $(u_{a}^{+})_{M}\in L^{r}({\mathbb{R}}^{N})$ for all $r\in[p_{s}^{*},\infty]$. Hence, we get $\phi=(u_{a}^{+})_{M}^{\sigma}\in L^{p_{s}^{*}}({\mathbb{R}}^{N})$. By definition of ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ given in (1.2), we conclude that $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$.
By Theorem 3.1, $u_{a}$ is a weak solution of (SP). Taking $\phi$ as a test function, we write
@@
Using ($\bf i{2}$), Remark 3.2, and ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, we have the following lower bound of the left hand side:
@@
Step 1: In this step, for $\sigma=p^{*}_{s}$, we show that the set $\{(u_{a}^{+})^{\sigma+p-1}:a\in(0,a_{1})\}$ is uniformly bounded on $L^{\frac{\sigma}{p}}({\mathbb{R}}^{N})$. From ($\bf\tilde{f{1}}$) and ($\bf f{1}$), for any $\epsilon>0$, there exists $C=C({\epsilon},a_{1})$ such that $\lvert f_{a}(u_{a})\rvert\leq C+\epsilon\lvert u_{a}\rvert^{p^{*}_{s}-1}$ for all $a\in(0,a_{1})$. We can also write the previous inequality as follows:
$\lvert f_{a}(u_{a})\rvert\leq C+\epsilon(u_{a}^{+}+u_{a}^{-})^{p^{*}_{s}-1}=C+\epsilon\left((u_{a}^{+})^{p^{*}_{s}-1}+(u_{a}^{-})^{p^{*}_{s}-1}\right).$ (e0)
Since $u_{a}^{+}$ and $u_{a}^{-}$ have disjoint supports, we estimate the right hand side of (3.8) using (e0) as follows:
$\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\leq C\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x+{\epsilon}\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x.$ (e1)
First we estimate the first integral on the right hand side of (e1). We use the continuous embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, the uniform boundedness of $\{u_{a}:a\in(0,a_{1})\}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ (Theorem 3.1), and Remark 3.2 to deduce that
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\leq C(N,s,p)\left\lVert g\right\rVert_{\infty}\left\lVert u_{a}^{+}\right\rVert_{s,p}^{p^{*}_{s}}\leq C(N,s,p,f,g,a_{1}),\quad\forall\,a\in(0,a_{1}).$ (e2)
In order to estimate the second integral on the right hand side of (e1), we use the Hölder’s inequality with the conjugate pair $(\frac{p^{*}_{s}}{p^{*}_{s}-p},\frac{p^{*}_{s}}{p})$ and we get
@@
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x\leq C(N,s,p,f,g,a_{1})\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}.$ (e3)
Combining (3.8), (e1), (e2) and (e3), we obtain
$\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\left(\frac{p^{*}_{s}+p-1}{p}\right)^{p}\left(C_{1}+C_{2}{\epsilon}\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\right),$ (3.9)
where $C_{1}=C_{1}(N,s,p,{\epsilon},f,g,a_{1})$ and $C_{2}=C_{2}(N,s,p,f,g,a_{1})$. Now we choose ${\epsilon}>0$ such that
@@
Using the above choice of $\epsilon>0$, the inequality (3.9) reduces to
@@
$\delta_{0}\leq\frac{\|u_{a}\|_{s,p}^{p}}{p}=\frac{1}{p}\int_{{\mathbb{R}}^{N}}g(x)|f_{a}(u_{a})u_{a}|\,{\rm d}x\leq C(f,a_{2})\|g\|_{1}(\|u_{a}\|_{\infty}+\|u_{a}\|_{\infty}^{p_{s}^{*}}).$ (3.16)
@@
Now we define $w_{n}=u_{n}-\tilde{u}$. Since $u_{n},\tilde{u}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{r}({\mathbb{R}}^{N})$ (Proposition 3.3 and Proposition 3.5), we have $w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty]$. Taking $\alpha=w_{n}(x)$ and $\beta=w_{n}(y)$ into the inequality (A) and following the same arguments in the proof of Proposition 3.3, where it is shown that $\phi=((u_{a}^{+})_{M})^{\sigma}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we get $\lvert w_{n}\rvert^{q-1}w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ for every $q\geq 1$.
For $p\geq 2$: We use [31, Lemma 2.3] to obtain the following lower bound of the right hand side of (3):
@@
where $C=C(N,s,p,f,a_{2})$ and $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Define $\theta:=\frac{p^{*}_{s}}{p\bar{r}^{\prime}}.$ Using the definition of $\bar{r}$ in (3.23) it follows that $\theta>1$. We consider two sequences $\{l_{j}\}$ and $\{m_{j}\}$ by the following recursive process:
$l_{0}=p_{s}^{*},\quad l_{j+1}=\theta l_{j}+\frac{p^{*}_{s}(p-1)}{p},\quad m_{j}=\frac{l_{j}}{\bar{r}^{\prime}}.$
@@
$\leq\left(\widetilde{C}m_{j}^{p}\right)^{\theta\bar{r}^{\prime}}\left\lVert g\right\rVert_{\infty}^{\theta}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta}$
$\leq\left(\widetilde{C}m_{j}^{p}\right)^{\theta\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta},$ (3.26)
for some $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. By Lemma A.2 we observe that $m_{j}\sim\frac{\theta^{j}}{\bar{r}^{\prime}}$ as $j\rightarrow\infty$. Thus iterating on (3) we obtain
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x$ $\leq\prod_{i=0}^{j-1}\left(\widetilde{C}m_{i}^{p}\right)^{\theta^{j-i}\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{0}}\,{\rm d}x\right)^{\theta^{j}}$
$\leq\left(\widetilde{C}\right)^{\sum\limits_{i=0}^{j-1}\theta^{j-i}\bar{r}^{\prime}}\theta^{p\bar{r}^{\prime}\sum\limits_{i=0}^{j-1}i\theta^{j-i}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\theta^{j}}.$ (3.27)
Notice that $S_{1}:=\sum\limits_{i=0}^{\infty}\theta^{-i}<\infty$ and $S_{2}:=\sum\limits_{i=0}^{\infty}i\theta^{-i}<\infty$. Take $C>\max\{1,\widetilde{C}\}$. Then for all $n\geq n_{1},j\in{\mathbb{N}}$ using (3) we deduce
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\leq C^{\theta^{j}S_{1}}\theta^{p\bar{r}^{\prime}\theta^{j}S_{2}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\theta^{j}}.$
@@
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C^{\frac{\theta^{j}}{l_{j}}S_{1}}\theta^{p\bar{r}^{\prime}\frac{\theta^{j}}{l_{j}}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\frac{\theta^{j}}{l_{j}}}\leq C^{\beta_{2}S_{1}}\theta^{p\bar{r}^{\prime}\beta_{2}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\beta_{1}},$
where $\beta_{1},\beta_{2}>0$ have been chosen such that for all $j\in{\mathbb{N}}$, $\beta_{1}<\frac{\theta^{j}}{l_{j}}<\beta_{2}$ (see Lemma A.2). Therefore, there exists $C>1$ such that for all $n\geq\max\{n_{1},n_{2}\}$ and $j\in{\mathbb{N}}$ large enough
@@
For $\frac{2N}{N+2s}<p<2$: For any $q\geq 1$, we take $\lvert w_{n}\rvert^{q-1}w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ as a test function in (3) and use [31, Lemma 2.4] to obtain
@@
where $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Define $\theta:=\frac{p^{*}_{s}}{2\bar{r}^{\prime}}.$ Using the definition of $\bar{r}$ in (3.30) we can verify that $\theta>1$. We consider two sequences $\{l_{j}\}$ and $\{m_{j}\}$ by the following recursive process:
$l_{0}=p_{s}^{*},\quad l_{j+1}=\theta l_{j}+\frac{p^{*}_{s}}{2},\quad m_{j}=\frac{l_{j}}{\bar{r}^{\prime}}.$
@@
$\leq\left(\widetilde{C}m_{j}^{2}\right)^{\theta\bar{r}^{\prime}}\left\lVert g\right\rVert_{\infty}^{\theta}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta}$
$\leq\left(\widetilde{C}m_{j}^{2}\right)^{\theta\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta},$ (3.32)
@@
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\leq C^{\theta^{j}S_{1}}\theta^{2\bar{r}^{\prime}\theta^{j}S_{2}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\theta^{j}},$
@@
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C^{\frac{\theta^{j}}{l_{j}}S_{1}}\theta^{2\bar{r}^{\prime}\frac{\theta^{j}}{l_{j}}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\frac{\theta^{j}}{l_{j}}}\leq C^{\alpha_{2}S_{1}}\theta^{2\bar{r}^{\prime}\alpha_{2}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\alpha_{1}},$
where $\alpha_{1},\alpha_{2}$ have been chosen such that for all $j\in{\mathbb{N}}$, $\alpha_{1}<\frac{\theta^{j}}{l_{j}}<\alpha_{2}$ (see Lemma A.2). Finally, we find a $C>1$ such that for all $n\geq\max\{n_{1},n_{3}\}$ and $j\in{\mathbb{N}}$ large enough
@@
$\Psi(\varrho)=\mathcal{H}^{N-2}(\mathbb{S}^{N-2})\int_{-1}^{1}\frac{(1-t^{2})^{\frac{N-3}{2}}}{(1-2t\varrho+\varrho^{2})^{\frac{N+sp}{2}}}\,\mathrm{d}t,$ (3.37)
where $\mathcal{H}^{N-2}$ is the Lebesgue measure of dimension $(N-2)$ and $\mathbb{S}^{N-2}$ is a unit sphere in $\mathbb{R}^{N-1}$.
@@
$|x-y|\geq|x|-|y|\geq|x|-\varepsilon\geq|x|-\frac{|x|}{2}=\frac{|x|}{2}.$
@@
where $C_{1}(\beta,\varepsilon)<0$. Since $a_{n}\rightarrow 0$, there exists $n_{1}\in{\mathbb{N}}$ such that $a_{n}B\leq-C_{1}(\beta,\varepsilon){\varepsilon}^{(\beta+1)(p-1)}$ and $u_{a_{n}}\geq 0$ a.e. in ${\mathbb{R}}^{N}$ (Proposition 3.8) for all $n\geq n_{1}$. Thus, for all $n\geq n_{1}$ using the assumption (3.38), the following holds weakly in $\overline{B_{R_{1}}}^{c}$: